OpenStudy (1645323):

Ms. Smith has decided to open a hot dog stand. She has found that her daily costs are approximated by the following equation: C(x) = x2 – 40x + 610. C(x) is the cost in dollars to sell x hot dogs. a) Find the number of hot dogs she must sell to minimize her cost. ________ b) Find that minimum cost. _______________ c) What is Ms. Smith’s beginning cost each day for the hot dog stand? __________________

OpenStudy (1645323):

@AravindG @skullpatrol @AnswerMyQuestions @greenlegodude57

OpenStudy (1645323):

@phi @OneRepublic @crabby_patty

OpenStudy (1645323):

is the beginning cost 610?

OpenStudy (anonymous):

Not sure. Sorry.

OpenStudy (phi):

The cost is the equation C(x) = x^2 – 40x + 610 that is the equation of a parabola with a U shape. Its vertex (at its bottom) is the minimum cost. do you know how to find the x value of the vertex of the equation y= a x^2 + bx + c ? it is when x= -b/(2a)